OpenStudy (mariooandresss):

Solve 4 over x plus 2 equals 2 over x for x and determine if the solution is extraneous or not.

OpenStudy (anonymous):

\[\frac{4}{x+2}=\frac{2}{x}\]??

OpenStudy (mariooandresss):

mhmm

OpenStudy (anonymous):

this is why mathematicians invented notation instead of using words

OpenStudy (anonymous):

if you like you can start by flipping both sides and solve \[\frac{x+2}{4}=\frac{x}{2}\]

OpenStudy (anonymous):

or just go right to \[4x=2(x+2)\]

OpenStudy (mariooandresss):

are w solving for 'x'

OpenStudy (mariooandresss):

we*

OpenStudy (anonymous):

\[\frac{a}{b}=\frac{c}{d}\iff ad=bc\] is always true yes you are solving for \(x\) there is not other variable there

OpenStudy (mariooandresss):

x=2

OpenStudy (anonymous):

yes that is correct if \(x=2\) then both sides of the equal sign are \(1\)

OpenStudy (mariooandresss):

correct, i don't know what the following step is though

OpenStudy (mariooandresss):

to determine if this is extraneous or not

OpenStudy (mariooandresss):

Is it extraneous? lets plug it in. \[\frac{4}{\color{red} {2}+2}=\frac{2}{\color{red} {2}}\]\[\frac{4}{4}=\frac{2}{2}\]So, yes, it is extraneous.

OpenStudy (mariooandresss):

thank you @satellite73